OpenStudy (vera_ewing):

Algera 2 Solving Polynomial Equations?!

OpenStudy (vera_ewing):

Welcome to Oakville Lake Amusement Park! As one of the new roller coaster engineers, you have been tasked with developing a roller coaster that will intertwine with existing Oakville Lake Amusement Park structures. For one of the more thrilling sections, the roller coaster will dive down in-between buildings, plummet underground, pop back up, and coast over a hill before shooting back underground. There must be three distinct points where the roller coaster crosses the x–axis. Precise measurements and attention to detail are very important.

OpenStudy (vera_ewing):

First, here is the existing map of current structures. It is important that the roller coaster does not go through the foundation of any of these structures. 1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___ 1. Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your roller coaster. Show all of your work. 2. Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences. 3. Solve for the y–intercept for your function, f(x), and then construct a rough graph of your roller coaster. If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (vera_ewing):

@amistre64 Please help me!

OpenStudy (amistre64):

ive actually work on this a few times before http://openstudy.com/study#/updates/54888264e4b079b941e803cc is one of the times

OpenStudy (amistre64):

how do we use the zeros of a function to construct a poly?

OpenStudy (vera_ewing):

Oh I'm not sure....

OpenStudy (amistre64):

im not here to fight with a time frame ....

OpenStudy (vera_ewing):

Ok so how do I do number 1?

OpenStudy (amistre64):

given a set of roots (a,b,c,...,k) we can create a polynomial that has those roots by making a product of the roots subtracted from x f(x) = (x-a)(x-b)(x-c)...(x-k)

OpenStudy (vera_ewing):

Ok so I know how to do #1, but I don't understand #2 and #3

OpenStudy (amistre64):

define for me fundamental Theorem: and Descartes` rule of signs:

OpenStudy (vera_ewing):

The fundamental theorem of a field of mathematics is the theorem considered central to that field. The naming of such a theorem is not necessarily based on how often it is used or the difficulty of its proofs. In mathematics, Descartes' rule of signs, first described by René Descartes in his work La Géométrie, is a technique for determining an upper bound on the number of positive or negative real roots of a polynomial. It is not a complete criterion, because it does not provide the exact number of positive or negative roots.

OpenStudy (vera_ewing):

@amistre64

OpenStudy (amistre64):

was just reviewing the post i did, we covered the solutions yes :) the fundamental thrm of algebra tells us that the degree of a polynomial defines how many roots it has. how many roots do we have?

OpenStudy (vera_ewing):

3?

OpenStudy (amistre64):

yeah, so we know that given the 3 different roots, the poly that we formed must satisfy the fundamental thrm by default: and that there are three distinct points where the roller coaster crosses the x–axis.

OpenStudy (amistre64):

gotta love the diamonds .... something about an encoding blurb in a copy paste

OpenStudy (amistre64):

in order to determine the rule of sign stuff, we will actually need the polynomial to play with

OpenStudy (vera_ewing):

Ok so now what? What would I say to prove to the construction foreman that my funtion matches the graph?

OpenStudy (vera_ewing):

(x-6) (x+2) x(x+2) -6(x+2) xx+2x -6x+2(-6) (x+7) (xx -4x -12) ^^^^^^^^ gotta wrap it for 'properness' x(xx -4x -12) +7(xx -4x -12) xxx -4xx -12x +7xx -4(7)x -12(7) xxx +3xx -12x -28x -84 xxx +3xx -40x -84

OpenStudy (vera_ewing):

You said that on the other question ^^

OpenStudy (amistre64):

and i believe i was right :) so f(x) = x^3 +3x^2 -40x -84 as far as the fundamental thrm goes, you will have to concult your material to see how it is better defined, it should basically say that 3 distinct roots form a polynomial of degree 3 by default.

OpenStudy (vera_ewing):

Ok so f(x) = x^3 +3x^2 -40x -84 is the polynomial for #1, and for #2 how do I explain?

OpenStudy (amistre64):

so descartes rule of sign is a little more involved, but its just counting really. count the number of times the operation changes from + to - in the poly (the start is always a zero count): f(x) = x^3 +3x^2 -40x -84 ^ ^ 0 1 by default we can only have 1 real positive root.

OpenStudy (vera_ewing):

So I can just say that ^ for #2?

OpenStudy (amistre64):

next we change x to -x and do the same thing f(-x) = (-x)^3 +3(-x)^2 -40(-x) -84 f(-x) = -x^3 +3x^2 +40x -84 ^ ^ ^ 0 1 2 so there are at most 2 negative real roots, or by some odd course of nature, there might be no real negative roots and 2 complex roots.

OpenStudy (vera_ewing):

Ok is that all for #2?

OpenStudy (amistre64):

what you say is up to you, im just trying to convey some information that might help you construct your own answers.

OpenStudy (amistre64):

all that is for descartes rule yes

OpenStudy (vera_ewing):

Ok so how about #3? how do i do that?

OpenStudy (amistre64):

how do we determine a y intercept?

OpenStudy (amistre64):

the graph and stuff i believe in on the other post i did

OpenStudy (vera_ewing):

I'm not really sure...It says to solve for the y–intercept for your function, f(x), and then construct a rough graph of your roller coaster. If your y–intercept is off the graph, give the coordinates of the y–intercept. I have no idea how to do that

OpenStudy (amistre64):

what is our function?

OpenStudy (vera_ewing):

f(x) = x^3 +3x^2 -40x -84

OpenStudy (amistre64):

now recall that the y axis has another name, the x=0 axis so to find the value of y intercept, we simply let x=0, what do we end up with?

OpenStudy (vera_ewing):

0?

OpenStudy (amistre64):

x^3 +3x^2 -40x -84, if x=0 this does not go to 0

OpenStudy (vera_ewing):

What does it go to?

OpenStudy (amistre64):

let x=0 and work the calculation ....

OpenStudy (vera_ewing):

Wait but I got 0? How did you not get 0?

OpenStudy (amistre64):

x^3 +3x^2 -40x -84 let x=0 0^3 +3(0)^2 -40(0) -84 0 +0 -0 -84 does not equal 0

OpenStudy (vera_ewing):

so it's -84?

OpenStudy (amistre64):

of course

OpenStudy (vera_ewing):

So is that the answer to #3 or is there more?

OpenStudy (amistre64):

make the rough graph of it ....

OpenStudy (vera_ewing):

and then is that it? that's it for the whole thing after i make the graph?

OpenStudy (amistre64):

other than organizing it into your own words ... yeah

OpenStudy (vera_ewing):

How would you say it?

OpenStudy (amistre64):

ive already said it :)

OpenStudy (vera_ewing):

No but for #2 it says explain. You already did that? @amistre64

OpenStudy (amistre64):

i explained it the best i could yes.

OpenStudy (vera_ewing):

Did you get my message about the graphing website?

OpenStudy (amistre64):

sketching a rough graph is simply makeing tics where the curve crosses the axises

OpenStudy (amistre64):

if you wanted to use a graphing site, yes you would want to use the equation we came up with

OpenStudy (vera_ewing):

Ok thank you so much! I really appreciate your help. Thanks for being so patient with me lol :)

OpenStudy (amistre64):

youre welcome :) good luck and all